OpenStudy (anonymous):

36/24 has how many halves

OpenStudy (saifoo.khan):

i think 3.

OpenStudy (anonymous):

I got 4 :\

OpenStudy (saifoo.khan):

how? :O

OpenStudy (ybarrap):

36/24 = x * 1/2 x = 2 * 36/24 = 36/12 = 3 Ans. 36/24 has 3 halves

OpenStudy (anonymous):

oh I was off by one!! I divided :|

OpenStudy (saifoo.khan):

Oh I see.